# LightStereo: Channel Boost Is All Your Need for Efficient 2D Cost Aggregation

### v1

@@ Abstract
We present LightStereo, a cutting-edge stereo-matching network crafted to accelerate the matching process. Departing from conventional methodologies that rely on aggregating computationally intensive 4D costs, LightStereo adopts the 3D cost volume as a lightweight alternative. While similar approaches have been explored previously, our breakthrough lies in enhancing performance through a dedicated focus on the channel dimension of the 3D cost volume, where the distribution of matching costs is encapsulated. Our exhaustive exploration has yielded plenty of strategies to amplify the capacity of the pivotal dimension, ensuring both precision and efficiency. We compare the proposed LightStereo with existing state-of-the-art methods across various benchmarks, which demonstrate its superior performance in speed, accuracy, and resource utilization. LightStereo achieves a competitive EPE metric in the SceneFlow datasets while demanding a minimum of only 22 GFLOPs, with an inference time of just 17 ms. Our comprehensive analysis reveals the effect of 2D cost aggregation for stereo matching, paving the way for real-world applications of efficient stereo systems. Code will be available at https://github.com/XiandaGuo/OpenStereo.
Figure 1: Performence vs. Time on SceneFlow(Left) and KITTI15(Right) datasets.
@@
11footnotetext: These authors contributed to the work equally.22footnotetext: Corresponding authors:long.chen@ia.ac.cn
1 Introduction
@@
Stereo matching is a pivotal task in computer vision, which aims to ascertain correspondences between points in stereo image pairs to compute depth information. This process underpins numerous applications, including autonomous driving, robotic navigation, and augmented reality. Despite substantial advancements, achieving real-time stereo matching without sacrificing accuracy or efficiency remains a formidable challenge, especially on resource-constrained platforms.
Recent developments gcnet ; psmnet2018 ; gwcnet2019 ; ganet2019 ; dsmnet2020 ; gu2020cascade ; leastereo ; wang2021pvstereo ; song2021adastereo ; Crestereo ; liu2022local ; acvnet ; nie2019multi ; guo2023openstereo have primarily focused on leveraging deep learning techniques for accurate stereo matching. Most methods are based on 3D CNN for cost aggregation. The pioneer is GCNet gcnet which jointly learns geometry and context by employing 3D CNN for cost aggregation. Due to the disparity dimension being restored in the constructed 4D cost volume, the network can perform exhaustively aggregation for accurate matching. Following works psmnet2018 ; gwcnet2019 also adopt this diagram and make notable advancements in performance. To mitigate the expenses in memory and computation of 3D cost aggregation, GANet ganet2019 proposes semi-global guided aggregation and local guided aggregation modules to replace 3D CNNs. LEAStereo leastereo utilizes a neural architecture search approach to automatically discover efficient and effective 3D cost aggregation architectures in stereo matching. CoEx bangunharcana2021coex employs a guided cost volume excitation approach to facilitate real-time stereo matching with enhanced accuracy. However, the use of 3D CNNs for cost aggregation in CoEx presents challenges for deployment on edge devices due to their computational demands. Another line of work is leveraging iterative refinement at multiple levels raftstereo ; xu2023iterative ; guo2023openstereo ; SelectiveStereo to refine disparity prediction on constructed all-pair 3D cost volume. Nevertheless, the overall runtime of these iterative-based methods is over 100ms on a custom GPU.
@@
There are also efforts wang2020fadnet ; wang2021fadnet++ ; shamsafar2022mobilestereonet ; xu2020aanet ; xu2023cgi focusing on lightweight design for stereo matching. AANet xu2020aanet constructs a 3D cost volume by correlating the left and right images and introduces intra-scale and cross-scale 2D cost aggregation modules to enhance the efficiency and accuracy of cost aggregation. MobilenetStereo-2D shamsafar2022mobilestereonet introduces MobileNet mijwil2023mobilenetv1 ; sandler2018mobilenetv2 blocks to reduce the cost, but the performance is far from satisfaction (EPE 1.14). Overall, these methods based on 2D cost aggregation perform poorly. Cost aggregation is critical to accuracy and efficiency in stereo matching yet existing methods require a compromise between accuracy and speed. However, existing methodologies often necessitate a trade-off between accuracy and processing speed. This paper poses the question: Is it possible to design a lightweight 2D encoder-decoder aggregation net to achieve precise disparity estimation?
@@
From the perspective of cost aggregation in stereo matching, focusing on the disparity channel dimension presents several advantages. Firstly, it allows for more direct modeling of the disparities between corresponding image points, which is crucial for accurate disparity estimation. By focusing on this dimension, it becomes possible to more effectively capture and assimilate the critical information needed for robust stereo matching. In this paper, we propose LightStereo as a positive solution. We explore 2D cost aggregation for stereo matching and leverage inverted residual blocks to enhance both accuracy and computational efficiency. The model is specifically designed to address the challenges of real-time stereo vision applications, focusing on reducing computational demands without compromising the quality of disparity estimation. Specifically, we utilize inverted residual blocks for 2D cost aggregation in stereo matching, which focuses on the disparity channel dimension of the 3D cost volume rather than the dimension of height and width. In inverted residual blocks, the expansion phase increases the channel dimension, allowing the network to learn richer features at a reduced computational cost before compressing them back. In addition, inspired by the effectiveness of large kernel and strip convolutions in image segmentation peng2017large ; hou2020strip ; guo2022segnext , we propose a Multi-Scale Convolutional Attention Module (MSCA) module for enhancing cost aggregation by extracting features from the left image. By leveraging multi-scale image features to excite the channel dimension of the cost volume, we utilize semantic information inherent in the images (such as object-level semantic details) to guide the cost aggregation process. When encountering discontinuities in disparity, the network halts propagation.
@@
  • We propose LightStereo, which achieves a competitive epe in the SceneFLow sceneflow datasets while demanding a minimum of only 22 Gflops with an inference time of 17 ms.
@@
  • We verify the effectiveness of LightStereo, achieving SOTA performance on the Sceneflow sceneflow and KITTI kitti2012 ; kitti2015 benchmark within the lightweight stereo-matching methods.
2 Related Work
Stereo matching, which predicts disparities (depth) from stereo images, can be classified into Conventional and deep-learning-based methods. Regardless of the method studied over the past few decades, researchers have constantly searched for the best trade-off between accuracy and speed to achieve better performance.
2.1 Conventional Stereo Matching
Stereo algorithms usually include the following four steps scharstein2002taxonomy : matching cost computation, cost aggregation, disparity computation/optimization, and disparity refinement. For correspondence problems in cost aggregation, the local method only uses the gray level, color, gradient, and other information of a certain neighborhood to calculate the matching cost, which has low computational complexity hosni2009local ; bleyer2011patchmatch ; hosni2012fast . Instead, Global approaches use pixel-based matching cost to search for disparity assignments that minimize the energy function over the entire stereo pair terzopoulos1986regularization and relative research mainly focuses on minimization procedure being used such as Graph Cut boykov2001fast , Markov Random Fields yamaguchi2014efficient and Dynamic Programming ohta1985stereo . Combined with the above two methods, SGM hirschmuller2005accurate still adopts the global framework, but uses a more efficient one-dimensional path aggregation method to replace the two-dimensional minimization algorithm. SGM hirschmuller2005accurate significantly improves the algorithm’s efficiency while being competitive with global algorithms in terms of accuracy.
2.2 Deep-learning-based Stereo Matching
To deal with the problem of specular surfaces, ambiguous regions, repetitive patterns, occlusions, and discontinuities, many modern studies use deep learning approaches to replace some or even all steps in the traditional matching process. These deep-learning-based stereo-matching algorithms can be categorized into two groups: accuracy-enhanced and real-time optimized.
@@ 2.2 Deep-learning-based Stereo Matching
Accuracy-Focused Stereo Matching. Many contemporary stereo-matching methods are dedicated to enhancing accuracy through various techniques and optimizations. Among these, 3D end-to-end networks have introduced significantly heightened precision in disparity estimation within stereo matching psmnet2018 ; gwcnet2019 ; leastereo ; acvnet ; xu2023iterative ; guo2023openstereo . PSMNet psmnet2018 adopts an architecture that leverages spatial pyramid pooling and 3D convolutional neural networks to learn disparity estimation from stereo images. GwcNet gwcnet2019 introduces a novel approach for constructing the cost volume in stereo matching using group-wise correlation. By dividing left and right features into groups along the channel dimension, correlation maps are computed within each group to generate multiple matching cost proposals. LEAStereo leastereo employs a hierarchical neural architecture search (NAS) framework to enhance deep stereo matching performance. IGEV xu2023iterative constructs a unified geometry encoding volume, integrating geometry, contextual cues, and local matching intricacies. OpenStereo guo2023openstereo conducted a comprehensive benchmark with a focus on practical applicability and introduced StereoBase, which further elevates the performance ceiling of stereo matching. To further optimize Stereo matching methods based on iterative optimization, Selective-IGEV SelectiveStereo proposes the Selective Recurrent Unit to help integrate disparity information across frequencies, minimizing loss during iterative steps. In addition, work  ganetADL improves the loss function by introducing ADL, an adaptive multi-modal cross-entropy loss, to guide network learning of diverse pixel distribution patterns.
@@
Real-Time-Focused Stereo Matching. To facilitate the real-time deployment of stereo-matching algorithms, a multitude of strategies have been proposed to optimize their practical application. StereoNet khamis2018stereonet uses color inputs to guide hierarchical refinement and can recover high-frequency details. DeepPrunerdeeppruner2019 proposes a PatchMatch module with learnable parameters to save memory and computation by gradually trimming the disparity space to be searched for each pixel. AnyNetanynet2019 has a 2D image convolutional network and a 3D cost tensor convolutional network. It computes disparity estimates at any time by continuously refining the disparity map resolution using the up-sampling method. FADNetwang2020fadnet leverages efficient 2D correlation layers with residual structures to perform multi-scale predictions. Based on the learned bilateral grid, BGNetbgnet designs an edge-preserving cost volume up-sampling module, allowing computationally expensive operations such as 3D convolution to be performed at low resolution. CoExbangunharcana2021coex uses image feature-guided weights and cost volume to excite 3D CNN to extract relevant geometric features. Building upon the foundation of StereoNet khamis2018stereonet , MobileStereoNetshamsafar2022mobilestereonet introduces two stereo vision models suitable for resource-constrained devices, achieving substantial reductions in both parameters and computational operations. However, existing stereo-matching methods still have considerable room for improvement in terms of computational efficiency and parameter scale.
@@
After carefully considering and testing various network structures, we propose a lightweight network, called LightStereo. This model innovatively explores 2D cost aggregation for stereo matching, leveraging advanced techniques to enhance both accuracy and computational efficiency.
Figure 2: The diagram above illustrates the architecture of LightStereo. MSCA refers to Multi-Scale Convolutional Attention module.
3 Method
@@
In the deployment of stereo matching on edge devices, constructing a 4D cost volume and using 3D CNNs for cost aggregation prove to be extremely inefficient. Our objective is to construct a 3D cost volume and utilize 2D CNNs, augmented with channel boost, for cost aggregation. This approach aims to maintain a balance between efficiency and accuracy, making it more suitable for real-time applications on resource-constrained devices. Our proposed LightStereo, as depicted in Figure 2, utilizes inverted residual blocks to aggregate disparity from a low-resolution cost volume. In this section, we first describe the design of channel-boosted inverted residual blocks for 2D cost aggregation (Section 3.1). Then, we introduce the Multi-Scale Convolutional Attention (MSCA) (Section 3.2). Finally, in Section 3.3, we present details of the network architecture of Lightstereo.
3.1 Inverted Residual Blocks for 2D Cost Aggregation
@@
Previous efforts, such as MobilenetStereo-2D shamsafar2022mobilestereonet , have incorporated MobileNet mijwil2023mobilenetv1 ; sandler2018mobilenetv2 blocks to decrease computational costs. Despite these efforts, the results have not met expectations, with an EPE of 1.14 on SceneFLow sceneflow still being reported. To address this issue, our study introduces inverted residual block to boost disparity estimation accuracy. As shown in Figure 3 (c), the inverted residual block sandler2018mobilenetv2 is a fundamental component in the design of a lightweight stereo matching network. Given the cost volume whose size is $H/4\times W/4\times Disp/4$, the key idea is first to boost the number of disparity channels, then apply depthwise convolution, and finally project the expanded features back to a lower-dimensional space, enhancing feature representation significantly. As shown in Figure 2, inverted residual blocks are utilized at resolutions of 1/4, 1/8, and 1/16, with each resolution corresponding to different blocks. For every block:
Initially, the cost volume $\mathbf{C}$ is first passed through a $1\times 1$ convolution to increase the number of disparity channel:
@@
where $\mathbf{W}_{\text{expand}}$ represents the weights of the expansion convolution, $\ast$ denotes the convolution operation, and $\sigma$ is the ReLU6 activation function. Next, the expanded feature map $\mathbf{y}$ undergoes a $3\times 3$ depthwise convolution, which operates independently on each disparity channel to capture spatial features:
@@
where $\mathbf{W}_{\text{depthwise}}$ are the weights of the depthwise convolution. Finally, the result $\mathbf{z}$ is then passed through another $1\times 1$ convolution to reduce the number of channels back to the original dimension:
@@
where $\mathbf{W}_{\text{project}}$ represents the weights of the projection convolution. If the input and output dimensions match, a skip connection is added to improve gradient flow and model performance:
@@
The inverted residual block’s structure significantly reduces computational complexity, making it ideal for resource-constrained environments. Our experimental results demonstrate the superiority of the inverted residual block design compared to regular CNN block (Figure 3 (a)), V1 block mijwil2023mobilenetv1 (Figure 3 (b)), and Vision Transformer (VIT) block liu2023efficientvit (Figure 3 (d)).
Figure 3: Comparison of different blocks for cost aggregation. DW refers to depthwise separable convolution. V1 Block represents the depthwise separable convolution mijwil2023mobilenetv1 . V2 Block represents an inverted residual block sandler2018mobilenetv2 . ViT Block refers to the block used in EfficientViT liu2023efficientvit .
3.2 Multi-Scale Convolutional Attention Module
@@
Inspired by the effectiveness of large kernel and strip convolutions in image segmentation peng2017large ; hou2020strip ; guo2022segnext , we propose a Multi-Scale Convolutional Attention Module (MSCA) module for enhancing cost aggregation by extracting features from the left image. The bottom-left corner of Figure 2 illustrates the architecture of the Multi-Scale Convolutional Attention (MSCA) module. This module is designed to capture and integrate features at multiple scales to enhance the feature representation for cost aggregation. The MSCA incorporates a series of depthwise separable convolutions with varying kernel sizes, specifically $1\times 1$, $7\times 1$, $1\times 7$, $11\times 1$, $1\times 1$ 11, $21\times 1$, and $1\times 21$. These convolutions capture both horizontal and vertical strip-like features, which are crucial for identifying elongated structures within the image. The reason we choose depth-wise strip convolutions is that they are lightweight. For example, by using a pair of $7\times 1$ and $1\times 7$ convolutions, we can effectively replace a standard $7\times 7$ convolution. Thus, strip convolutions serve as a complement to grid convolutions and aid in extracting strip-like features of left images. Utilizing multi-scale image features, MSCA enhances the channel dimension of the cost volume. MSCA incorporates semantic information embedded within the images to direct the cost aggregation process. The network is designed to cease propagation upon detecting disparities that are discontinuous.
@@
Given a stereo image input of size $H\times W\times 3$, the method obtains three scales of feature maps at 1/4, 1/8, and 1/16 of the original resolution, respectively. Then, these multi-scale feature maps are then processed by the MSCA module to further extract horizontal and vertical strip-like features. The outputs of MSCA are then concatenated to form a comprehensive multi-scale feature representation. This aggregated feature is subsequently processed by a $1\times 1$ convolution, which acts as a channel mixer. The $1\times 1$ convolution blends the multi-scale features and recalibrates the feature channels, enhancing the network’s ability to focus on relevant information across different scales. After the MSCA processing, the final output is combined with the aggregated cost by multiplication, enhancing the precision of cost aggregation.
3.3 Network Architecture of LightStereo

### v3

# LightStereo: Channel Boost Is All You Need for Efficient 2D Cost Aggregation

@@ Abstract
We present LightStereo, a cutting-edge stereo-matching network crafted to accelerate the matching process. Departing from conventional methodologies that rely on aggregating computationally intensive 4D costs, LightStereo adopts the 3D cost volume as a lightweight alternative. While similar approaches have been explored previously, our breakthrough lies in enhancing performance through a dedicated focus on the channel dimension of the 3D cost volume, where the distribution of matching costs is encapsulated. Our exhaustive exploration has yielded plenty of strategies to amplify the capacity of the pivotal dimension, ensuring both precision and efficiency. We compare the proposed LightStereo with existing state-of-the-art methods across various benchmarks, which demonstrate its superior performance in speed, accuracy, and resource utilization. LightStereo achieves a competitive EPE metric in the SceneFlow datasets while demanding a minimum of only 22 GFLOPs and 17 ms of runtime, and ranks 1st on KITTI 2015 among real-time models. Our comprehensive analysis reveals the effect of 2D cost aggregation for stereo matching, paving the way for real-world applications of efficient stereo systems. Code is available at https://github.com/XiandaGuo/OpenStereo.
@@
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding Author
I Introduction
@@
Stereo matching is a pivotal task in computer vision, which aims to ascertain correspondences between points in stereo image pairs to retrieve depth information. This process underpins numerous applications, including autonomous driving, robotic navigation, and augmented reality. Despite substantial advancements, achieving real-time stereo matching without sacrificing accuracy or efficiency remains a formidable challenge, especially on embedded platforms.
Most developments [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] have primarily focused on leveraging deep learning techniques for accurate stereo matching. Most methods are based on 3D CNN for cost aggregation: because of the disparity dimension being modeled in the constructed 4D cost volume, the network can perform exhaustively aggregation for accurate matching, yet at the expense of high memory consumption and runtime. Nevertheless, the overall runtime of these iterative-based methods is over 100ms on a custom GPU.
Figure 1: Performence vs. Time trade-off. Results on SceneFlow (Left) and KITTI15 (Right) datasets.
Figure 2: The diagram above illustrates the architecture of LightStereo.
@@
There are also efforts [15, 16, 17, 18, 19] focusing on lightweight design for stereo matching. AANet [18] constructs a 3D cost volume by correlating the left and right images and introduces intra-scale and cross-scale 2D cost aggregation modules to enhance the efficiency and accuracy of cost aggregation. MobilenetStereo-2D [17] introduces MobileNet [20, 21] blocks to reduce the cost, but the performance is far from satisfaction. Overall, these methods based on 2D cost aggregation perform poorly. Cost aggregation is critical to accuracy and efficiency in stereo matching yet existing methods require a compromise between accuracy and speed. However, existing methodologies often necessitate a trade-off between accuracy and processing speed. This paper poses the question: Is it possible to design a lightweight 2D encoder-decoder aggregation net to achieve precise disparity estimation?
@@
From the perspective of cost aggregation in stereo matching, focusing on the disparity channel dimension presents several advantages. Firstly, it allows for more direct modeling of the disparities between corresponding image points, which is crucial for accurate disparity estimation. By focusing on this dimension, it becomes possible to more effectively capture and assimilate the critical information needed for robust stereo matching. In this paper, we propose LightStereo as a positive solution. We explore 2D cost aggregation for stereo matching and leverage inverted residual blocks to enhance both accuracy and computational efficiency. The model is specifically designed to address the challenges of real-time stereo vision applications, focusing on reducing computational demands without compromising the quality of disparity estimation. Specifically, we utilize inverted residual blocks for 2D cost aggregation in stereo matching, which focuses on the disparity channel dimension of the 3D cost volume rather than the dimension of height and width. In inverted residual blocks, the expansion phase increases the channel dimension, allowing the network to learn richer features at a reduced computational cost before compressing them back. In addition, inspired by the effectiveness of large kernel and strip convolutions in image segmentation [22, 23, 24], we propose a Multi-Scale Convolutional Attention Module (MSCA) module for enhancing cost aggregation by extracting features from the left image. By leveraging multi-scale image features to excite the channel dimension of the cost volume, we utilize semantic information inherent in the images (such as object-level semantic details) to guide the cost aggregation process. When encountering discontinuities in disparity, the network halts propagation.
@@
  • We propose LightStereo, which achieves a competitive End-Point-Error (EPE) in the SceneFLow [25] datasets while demanding a minimum of only 22 Gflops with an inference time of 17 ms.
@@
  • We verify the effectiveness of LightStereo, achieving state-of-the-art performance on the Sceneflow [25] and KITTI [26, 27] benchmark within the lightweight stereo-matching methods.
II Related Work
We review the literature concerning deep learning-based stereo matching, identifying two categories of methods. We refer the reader to [28, 29] for a broader overview.
@@ II Related Work
Accuracy-Focused Stereo Matching. Many contemporary stereo-matching methods are dedicated to enhancing accuracy through various techniques and optimizations. Among these, 3D end-to-end networks have introduced significantly heightened precision in disparity estimation within stereo matching [2, 3, 7, 12, 30, 14]. PSMNet [2] adopts an architecture that leverages spatial pyramid pooling and 3D convolutional neural networks to learn disparity estimation from stereo images. GwcNet [3] introduces a novel approach for constructing the cost volume in stereo matching using group-wise correlation. By dividing left and right features into groups along the channel dimension, correlation maps are computed within each group to generate multiple matching cost proposals. LEAStereo [7] employs a hierarchical neural architecture search (NAS) framework to enhance deep stereo matching performance. IGEV [30] constructs a unified geometry encoding volume, integrating geometry, contextual cues, and local matching intricacies. OpenStereo [14] conducted a comprehensive benchmark with a focus on practical applicability and introduced StereoBase, which further elevates the performance ceiling of stereo matching. Selective-IGEV [31] proposes the Selective Recurrent Unit to help integrate disparity information across frequencies, minimizing loss during iterative steps. In addition, work  [32] improves the loss function by introducing ADL, an adaptive multi-modal cross-entropy loss, to guide network learning of diverse pixel distribution patterns.
@@
Real-Time-Focused Stereo Matching. A multitude of models have been proposed to target practical applications. StereoNet [33] uses color inputs to guide hierarchical refinement and can recover high-frequency details. DeepPruner[34] proposes a PatchMatch module with learnable parameters to save memory and computation by gradually trimming the disparity space to be searched for each pixel. AnyNet[35] has a 2D image convolutional network and a tiny 3D cost aggregation network. FADNet[15] leverages efficient 2D correlation layers with residual structures to perform multi-scale predictions. Based on the learned bilateral grid, BGNet[36] designs an edge-preserving cost volume up-sampling module, allowing computationally expensive operations such as 3D convolution to be performed at low resolution. MADNet series [37, 38] relies on fully 2D modules, maximizing speed at the expense of accuracy. CoEx[39] uses image feature-guided weights and cost volume to excite 3D CNN to extract relevant geometric features. Building upon the foundation of StereoNet [33], MobileStereoNet[17] introduces two models suitable for resource-constrained devices.
@@
However, these methods still have considerable room for improvement in terms of computational efficiency and parameter scale. After carefully considering and testing various network structures, we propose a lightweight network, called LightStereo. This model innovatively explores 2D cost aggregation for stereo matching, leveraging advanced techniques to enhance both accuracy and computational efficiency.
Figure 3: Comparison of different blocks used for cost aggregation – DW refers to depthwise separable convolution; V1 Block represents the depthwise separable convolution [20]; V2 Block represents an inverted residual block [21]; ViT Block refers to the block used in EfficientViT [40].
III Method
@@
In the deployment of stereo matching on edge devices, constructing a 4D cost volume and using 3D CNNs for cost aggregation prove to be extremely inefficient. Our objective is to construct a 3D cost volume, where single scores are stored along the channel dimension, and utilize 2D CNNs augmented with channel boost for cost aggregation to balance efficiency and accuracy. We are now going to introduce LightStereo, whose overview is shown in Figure 2, starting with its main components – the Inverted Residual Blocks [21] and the new Multi-Scale Convolutional Attention – and concluding with the overall architectural design.
III-A Inverted Residual Blocks for 2D Cost Aggregation
@@
Previous efforts, such as MobilenetStereo-2D [17], have incorporated MobileNet [20, 21] blocks to decrease computational costs. Despite these efforts, the results have not met expectations, with an EPE of 1.14 on SceneFLow [25] still being reported. To address this issue, our study introduces inverted residual blocks to boost disparity estimation accuracy, as shown in Figure 3 (c). Given the cost volume in Figure 2 whose size is $\frac{H}{4}\times\frac{W}{4}\times\frac{Disp}{4}$, the key idea is first to boost the number of disparity channels, then apply depthwise convolution, and finally project the expanded features back to a lower-dimensional space, enhancing feature representation significantly. Inverted residual blocks are utilized at resolutions of $\frac{1}{4}$, $\frac{1}{8}$, and $\frac{1}{16}$, each corresponding to different blocks: initially, the cost volume $\mathbf{C}$ is first passed through a $1\times 1$ convolution to increase the number of disparity channel:
@@
where $\mathbf{W}_{\text{expand}}$ represents the weights of the expansion convolution, $\ast$ denotes the convolution operation, and $\sigma$ is the ReLU6 activation function. Next, the expanded feature map $\mathbf{y}$ undergoes a $3\times 3$ depthwise convolution, which operates independently on each channel to capture spatial features:
@@
where $\mathbf{W}_{\text{depthwise}}$ are the weights of the depthwise convolution. Finally, the result $\mathbf{z}$ is then passed through another $1\times 1$ convolution to restore the original number of channels:
@@
where $\mathbf{W}_{\text{project}}$ represents the weights of the projection convolution. If the input and output dimensions match, a skip connection is added to improve gradient flow:
@@
The structure of the inverted residual block significantly reduces computational complexity, making it ideal for resource-constrained environments. Our experimental results will demonstrate its superiority over regular CNN block, V1 block [20], and Vision Transformer (ViT) block [40] – depicted in Figure 3 (a), (b) and (d) respectively.
Figure 4: Multi-Scale Convolutional Attention (MSCA).
III-B Multi-Scale Convolutional Attention Module
@@
Inspired by the effectiveness of large kernel and strip convolutions in image segmentation [22, 23, 24], we propose a Multi-Scale Convolutional Attention (MSCA) module for enhancing cost aggregation by extracting features from the left image. Figure 4 illustrates the architecture of such a module, designed to capture and integrate features at multiple scales and enhance their representation for cost aggregation. The MSCA incorporates a series of depthwise separable convolutions with varying kernel sizes, specifically $1\times 1$, $7\times 1$, $1\times 7$, $11\times 1$, $1\times 1$ 11, $21\times 1$, and $1\times 21$, capturing both horizontal and vertical strip-like features, which are crucial for identifying elongated structures within the image. These depth-wise strip convolutions are lightweight and can replace, for instance, a standard $7\times 7$ convolution with $7\times 1$ and $1\times 7$ convolutions. Thus, strip convolutions serve as a complement to grid convolutions and aid in extracting strip-like features of left images. Utilizing multi-scale image features, MSCA enhances the channel dimension of the cost volume, by incorporating semantic information embedded within the images to guide the aggregation process. The network is designed to cease propagation upon detecting disparity discontinuities.
@@
Given a stereo image of size $H\times W\times 3$, feature maps at $\frac{1}{4}$, $\frac{1}{8}$, and $\frac{1}{16}$ of the original resolution, and processed by the MSCA module to further extract horizontal and vertical strip-like features. The outputs are then concatenated to form a comprehensive multi-scale feature representation, subsequently processed by a $1\times 1$ convolution, which acts as a channel mixer. The $1\times 1$ convolution blends the multi-scale features and recalibrates the feature channels, allowing the network to focus on relevant information across different scales. The final output is combined with the aggregated cost by multiplication, enhancing the effect of cost aggregation.
III-C LightStereo – Network Architecture